Mercury Airlines’ marginal revenue and demand curves cross the marginal cost curve at
quantities of 3,000 and 6,000 seats a week, respectively. All other data remain the same.
a. Calculate the profit under policies of (i) uniform pricing, and
(ii) Complete price discrimination.
b. Suppose that Mercury implements complete price discrimination. Explain why it
should sell up to the quantity where the buyer’s marginal benefit equals Mercury’s
marginal cost.
1
Expert's answer
2018-03-12T16:11:07-0400
(i) Under the conditions of uniform pricing, the profit-maximizing quantity corresponds to MR=MC or Q=3000. The price p on the demand curve must be halfway between 800 and 4000.
Therefore, p = 2400, and profit equals = (2400 – 800) x 3000 = 4.8 million dirhams/week.

(ii) Under the condition of complete price discrimination, the seller should sell the quantity where MB = MC. Quantity is Q = 6000. Therefore, total revenue is the area under the demand curve up to Q = 6000, therefore: TR = [(4000 + 800)/2] x 6000 = 14.4 million dirhams / week.
Total cost will be, TC = 800 x 6000 = 4.8 million dirhams, therefore the profit will be: 14.4 – 4.8 = 9.6 million dirhams/week.

(b) Mercury Airlines maximizes own profit by producing the quantity where the buyer’s marginal benefit equals Mercury’s marginal cost. If it sold a larger quantity, so that the marginal benefit is smaller than the marginal cost, then its profit would be lower. If it sold a smaller quantity, it could grow profit by selling more.